### v2

Abstract
Self-supervised learning methods based on data augmentations, such as SimCLR, BYOL, or DINO, allow obtaining semantically meaningful representations of image datasets and are widely used prior to supervised fine-tuning. A recent self-supervised learning method, -SimCNE, uses contrastive learning to directly train a 2D representation suitable for visualisation. When applied to natural image datasets, -SimCNE yields 2D visualisations with semantically meaningful clusters. In this work, we used -SimCNE to visualise medical image datasets, including examples from dermatology, histology, and blood microscopy. We found that increasing the set of data augmentations to include arbitrary rotations improved the results in terms of class separability, compared to data augmentations used for natural images. Our 2D representations show medically relevant structures and can be used to aid data exploration and annotation, improving on common approaches for data visualisation.
keywords:
Self-supervised learning, augmentations, contrastive learning, data visualisation1 Introduction
Medical image datasets have been quickly growing in size and complexity [Litjens etĀ al. (2017), Topol (2019), Zhou etĀ al. (2021)]. Whereas medical professionals can analyse, annotate, and classify individual images, tasks involving large batches of images, ranging from data curation and quality control to exploratory analysis, remain challenging.
Self-supervised learning (SSL) has recently emerged in computer vision as the dominant paradigm for learning image representations suitable for downstream tasks [Balestriero etĀ al. (2023)], and it has increasingly been adopted in medical imaging [Huang etĀ al. (2023)]. In contrastive learning methods, such as SimCLR [Chen etĀ al. (2020)], BYOL (Grill etĀ al., 2020), or DINO Caron etĀ al. (2021), data augmentation is used to generate different views of each image, and a deep network is trained to keep these views close together in the representation space. However, the learned representations are typically high-dimensional.
Recently, Böhm et al. (2023) suggested a self-supervised contrastive method, called -SimCNE, for 2D visualisation of image datasets. Using natural image datasets, the authors demonstrated that -SimCNE obtains semantically meaningful visualisations, representing rich cluster structure and highlighting artefacts in the data. Their methods clearly outperformed existing 2D embedding methods like -SNE Van der Maaten and Hinton (2008) and UMAP McInnes et al. (2020) for natural image data.
[…]
2 Related work
Contrastive learning methods have been widely applied to medical image datasets (for a review, see Huang etĀ al., 2023) but usually as pre-training for downstream tasks such as classification or segmentation. Some recent works visualised high-dimensional SSL representations; e.g. Cisternino etĀ al. (2023) used UMAP of DINO to visualise histopathology data. In contrast, our focus is on self-supervised visualisations trained end-to-end.
Contrastive learning relies on data augmentations to create several views of each image, and the choice of data augmentations plays a crucial role in methodsā success Tian etĀ al. (2020). A large number of works explored data augmentations for medical images in a supervised setting (reviewed by Chlap etĀ al., 2021; Goceri, 2023). In the self-supervised context, vanĀ der Sluijs etĀ al. (2023) studied the effect of augmentations on the representation of X-ray images. For histopathology images, Kang etĀ al. (2023) advocated for using rotations and vertical flips, as well as staining-informed color transformations, while some other works used neighbouring patches as positive pairs Li etĀ al. (2021); Wang etĀ al. (2021).
3 Background: SimCLR and -SimCNE
SimCLR (Chen etĀ al., 2020) produces two augmentations for each image in a given mini-batch of size , resulting in augmented images. Each pair of augmentations forms a so-called positive pair, whereas all other possible pairs in the mini-batch form negative pairs. The model is trained to maximise the similarity between the positive pair elements while simultaneously minimising the similarity between the negative pair elements.
An augmented image is passed through a ResNet He etĀ al. (2016) backbone to give the latent representation , which is then passed through a fully-connected projection head with one hidden layer to yield the final output . SimCLR employs the InfoNCE loss function vanĀ den Oord etĀ al. (2019), which for one positive pair can be written as
where is the cosine similarity and is a hyperparameter that was set to in Chen etĀ al. (2020). Even though the loss function operates on (typically 128-dimensional), for downstream tasks, SimCLR uses the representations Bordes etĀ al. (2022), typically at least 512-dimensional.
The idea of -SimCNE Böhm et al. (2023) is to make the network output () two-dimensional so that it is directly suitable for data visualization. It does not make sense to apply the cosine similarity to this representation in as it would effectively normalise the embeddings to lie on a one-dimensional circle. -SimCNE replaces the exponential of the scaled cosine similarity with the Cauchy similarity used in -SNE (Van der Maaten and Hinton, 2008): . The resulting loss function is
Böhm et al. (2023) found that directly optimizing this loss is difficult, and suggested a three-stage process. The first stage (1000 epochs) used a 128-dimensional output which was then replaced with a 2D output and fine-tuned in the subsequent two stages (500 epochs).
For their experiments on CIFAR datasets, the authors used a ResNet18 with a modified first layer kernel size of , and a projection head with hidden layer size ofĀ 1024 (FigureĀ 1a).
4 Experimental setup
| Dataset | Image dim. | Sample size | Classes | Reference |
| Leukemia | Matek etĀ al. (2019a) | |||
| Bloodmnist | Yang etĀ al. (2023) | |||
| Dermamnist | Yang etĀ al. (2023) | |||
| Pathmnist | Yang etĀ al. (2023) | |||
| PCam16 | Veeling etĀ al. (2018) |
Datasets
We used five publicly available medical image datasets with sample sizes ranging from 10ā000 to over 300ā000 (TableĀ 1). Three datasets we took from the Medmnistv2 collection (Yang etĀ al., 2023), all consisting of RGB images. Dermamnist is based on the HAM10000 dataset Tschandl etĀ al. (2018), a collection of multi-source dermatoscopic images of common pigmented skin lesions. The images are categorised into 7 classes, which we reduced to binary labels: melanocytic nevi and other skin conditions. Bloodmnist is based on a dataset of microscopy images of individual blood cells from healthy donors Acevedo etĀ al. (2020), categorised into classes corresponding to cell types. Pathmnist is based on a dataset of non-overlapping patches from colorectal cancer histology slides Kather etĀ al. (2019), categorized into classes corresponding to tissue types. The Leukemia dataset Matek etĀ al. (2019b) contains microscopy images of white blood cells taken from patients, some of which were diagnosed with acute myeloid leukemia. We resized images to and merged 9 rare classes ( cells) into one, obtaining 7 classes.
The Patch Camelyon16 (PCam16) dataset Veeling etĀ al. (2018), adapted from the Camelyon16 challenge Bejnordi etĀ al. (2017), consists of patches from breast cancer histology slides. There are two classes: metastases and non-metastases. A patch was classified as metastases if there was any amount of tumor tissue in its central region.
Augmentations
Böhm et al. (2023) worked with natural images and used the same data augmentations as Chen et al. (2020): cropping, horizontal flipping, color jittering, and grayscaling (Figure 1b). Here we used all of these augmentations with the same hyperparameters and probabilities (see Table S1 for ablations). We reasoned that the semantics of microscopy or pathology images should be invariant to arbitrary rotations and arbitrary flips (Kang et al., 2023). For that reason we considered two additional sets of augmentations: (i) vertical flips and arbitrary 90° rotations; (ii) rotations by any arbitrary angle (Figure 1c). In each case, all possible rotations were equally likely. When rotating an image by an angle that is not a multiple of 90°, the corners need to be filled in (Figure 1c, right). For this we used the average color of all border pixels across all images in a given dataset. This color was dataset specific, but the same for all images in a dataset.
Architecture and training
We used the original -SimCNE implementation Böhm et al. (2023) with default parameters unless stated otherwise. For PCam16, we used the unmodified ResNet18 He et al. (2016) without the fully-connected layer. All networks were trained from scratch on an NVIDIA RTX A6000 GPU with the batch size of 1024, except for PCam16 where we had to reduce the batch size to 512 to fit it into GPU memory.
Baselines
For comparison, we applied -SNE to images in pixel space, in pretrained ResNet representation, and in SimCLR representation. The SimCLR models had the same architecture as -SimCNE models but with 128D output and were trained with SimCLR loss (Eq.Ā 1) for 1000 epochs. We then applied -SNE to the 512-dimensional SimCLR representation before the projector head. We took ImageNet-pretrained ResNet18 and ResNet152 models from the PyTorch library Paszke etĀ al. (2019). To pass our images through these networks, we resized all images to , center cropped to , and normalized (following He etĀ al., 2016). The resulting representations had 512 and 2048 dimensions respectively. We used openTSNE PoliÄar etĀ al. (2019) with default settings to reduce to 2D. When doing -SNE of the PCam16 data in pixel space, we first performed principal component analysis and only used the first PCs as input to -SNE.
Evaluation
We used two metrics to evaluate the quality of 2D embeddings, with classification and clustering being two possible downstream tasks: NN classification accuracy Pedregosa etĀ al. (2011) with and a 9:1 training/test split, and silhouette score (Rousseeuw, 1987). For a single point , the silhouette score is defined as where is the average distance between and points from the same class, and is the average distance between and points from the closest other class. The silhouette score of the entire embedding is the average across all points. These two measures are complementary: The NN accuracy measures how well the classes are separated from each other, while the silhouette score measures how far they are separated from each other.
5 Results
| Method | Dataset | |||||
| Leukemia | Bloodmnist | Dermamnist | Pathmnist | PCam16 | ||
| -SimCNE | def. augm. | |||||
| + 90° rot. | ||||||
| + rand. rot. | ||||||
| -SNEĀ of SimCLR | def. augm. | |||||
| + 90° rot. | ||||||
| + rand. rot. | ||||||
| -SNE | pixel space | |||||
| ResNet18 | ||||||
| ResNet152 | ||||||
| Method | Dataset | |||||
| Leukemia | Bloodmnist | Dermamnist | Pathmnist | PCam16 | ||
| -SimCNE | def. augm. | |||||
| + 90° rot. | ||||||
| + rand. rot. | ||||||
| -SNEĀ of SimCLR | def. augm. | |||||
| + 90° rot. | ||||||
| + rand. rot. | ||||||
| -SNE | pixel space | |||||
| ResNet18 | ||||||
| ResNet152 | ||||||
In this study, we asked (i) how the contrastive visualisation method -SimCNE (Böhm et al., 2023) could be applied to medical image datasets, and (ii) if the set of data augmentations could be enriched compared to what is typically used on natural images.
We considered the Leukemia dataset as our first example (FigureĀ 2). Naive application of -SNE to the raw images in pixel space resulted in an embedding with little class separation and low NN accuracy of 67.4% (FigureĀ 2a). Passing all images through an ImageNet-pretrained ResNet and then embedding them with -SNE improved the NN accuracy to 82.2% but visually the classes were still separated poorly (FigureĀ 2b). Training -SimCNE with default data augmentations gave embeddings with 86.7% NN accuracy (TableĀ 2) and much better visual class separation (FigureĀ 2c and TableĀ 3). This shows that -SimCNE can produce meaningful 2D visualizations of medical image datasets.
We reasoned that the set of data augmentations could be enriched to include 90° rotations and flips because the semantics of blood microscopy images is rotationally invariant. When training -SimCNE with this set of data augmentations, the NN accuracy increased to 94.0%. Additionally including all possible rotations by an arbitrary angle as data augmentations yielded the highest NN accuracy (95.2%) and the highest silhouette score (0.24), indicating that domain-specific augmentations can further improve -SimCNE embeddings.
Across the five datasets considered in this study, we saw three different outcomes. On microscropy datasets (Leukemia and Bloodmnist), -SimCNE with random rotations performed the best: it had by far the best silhouette score (TableĀ 3) and visually the most separated classes (FigureĀ 3a,b). SimCLR followed by -SNE has also benefited from rotational augmentations. Compared to -SimCNE, it had slightly higher NN accuracies (FigureĀ 2), but much lower silhouette scores.
On pathology datasets (Pathmnist and PCam16), -SimCNE with 90° rotations performed the best. On Pathmnist, it had the highest silhouette score (Figure 4). On PCam16, -SimCNE showed clearer structures compared to SimCLR + -SNE, but this difference was not captured by the silhouette scores which on this dataset were all close to zero (Table 3). This is because it only had two classes, whereas -SimCNE separated images not only by class but also by tissue types (Figure 5); this led to large within-class distances and hence misleadingly low silhouette scores.
Finally, on the dermatology dataset (Dermamnist), performance of all methods was similarly poor: SimCLR and -SimCNE resulted in embeddings not very different from -SNE in pixel space (FigureĀ 3c).
As a control experiment, we applied -SimCNE with 90° rotations and vertical flips to the CIFAR-10 dataset Krizhevsky et al. (2009). It decreased the NN accuracy from 89% to 76%. This confirms that rotation augmentations are not helpful for natural images because they are not invariant to rotations, unlike microscopy and pathology images.
In the pathology datasets, -SimCNE showed meaningful subclass structure. For example, in Pathmnist, the debris class separated into three clearly distinct subsets (FiguresĀ 4), one of which had markedly different staining colour. In the PCam16 dataset, the embedding clearly split patches with and without metastasis, based on the density of chromatin and variation in the size of the cells. The difference in visual appearance (different shades of violet) between top-right and bottom-left likely reveals a technical artefact resulting from different staining durations.
6 Discussion
In this paper, we showed that -SimCNE (Böhm et al., 2023) can be successfully applied to medical image datasets, yielding semantically meaningful visualisations, and benefits from rotational data augmentations, leveraging rotational invariance of microscropy images. In agreement with Böhm et al. (2023), -SimCNE performed better than SimCLR + -SNE combination. Even though SimCLR tended to have slightly higher NN accuracy, the silhouette score was typically much lower: -SimCNE achieved visually much stronger cluster separation, which is useful for practical visualisations. Furthermore, parametric nature of -SimCNE allows to embed new (out-of-sample) images into an existing embedding.
We found that blood microscopy datasets benefited the most from random rotations, while pathology datasets showed the best results with 90° rotations and flips. We believe it is because in blood microscopy images, the semantically meaningful part is always in the center (Figure 3a,b) and so the corners of the image may not be important. In contrast, in histopathology images, the edges of the image may contain relevant information, which may get rotated out of the image and replaced by solid-color triangles (Figure 1c). One of the datasets, Dermamnist, exhibited poor results with all analysis methods. This may be because in this dataset the images are too small to convey biomedically relevant information, or because the sample size was insufficient (Table 1).
In conclusion, we argue that -SimCNE is a promising tool for visualisation of medical image datasets. It can be useful for quality control, highlighting artefacts and problems in the data. It can also create a 2D map of cell types, tissue types, or medical conditions, which can be useful not only for clinical purposes but also education and research, potentially combined with an interactive image exploration tool. In the future, it may be interesting to extend -SimCNE to learn representations invariant to technical (e.g. staining) artefacts.
We thank Christian Schürch for discussion on histopathology data. This work was supported by the German Science Foundation (Excellence Cluster 2064 āMachine Learning ā New Perspectives for Scienceā, project number 390727645), the Hertie Foundation, and the Cyber Valley Research Fund (D.30.28739). The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Jan Niklas Bƶhm. Philipp Berens is a member of the Else Krƶner Medical Scientist Kolleg āClinbrAIn: Artificial Intelligence for Clinical Brain Researchā.
Appendix
| Augmentations | Leukemia | BloodMNIST | PathMNIST | |||
| NN acc. | Silhouette | NN acc. | Silhouette | NN acc. | Silhouette | |
| All | ||||||
| No crops | ||||||
| No color jitter | ||||||
| No grayscaling | ||||||